OpenStudy (anonymous):

write the equation for the (ellipse) conic section if the cone is centered at (0,0), the major axis 10 units, and the minor axis is 6 units

OpenStudy (tkhunny):

One piece at a time. It's an ellipse. It looks like this: \(\dfrac{(x-h)^{2}}{} + \dfrac{(y-k)^{2}}{} = 1\) Right?

OpenStudy (anonymous):

so it's the same as a circle?